OpenStudy (anonymous):

if f(x)=√ 2x+5, find the value of f^-1(5) help me pleaaasee :D

OpenStudy (anonymous):

f^-1(5) means f(x) = 5

OpenStudy (anonymous):

you're given f(x). Just find x

OpenStudy (anonymous):

doesnt f^-1 mean the inverse?

OpenStudy (anonymous):

so like the inverse of f(x) when f(x)=5? :/

OpenStudy (jdoe0001):

\(\bf f(x)={\color{red}{ y}}=\sqrt{2{\color{blue}{ x}}+5}\qquad inverse\implies {\color{blue}{ x}}=\sqrt{2{\color{red}{ y}}+5}\) solve for "y" notice that to find the inverse, all we do is swap about the variables, then solve for "y"

OpenStudy (anonymous):

ok, but what does the (5) part do? if anything?

OpenStudy (jdoe0001):

well, first off, find the inverse

OpenStudy (jdoe0001):

then once you find that, you'd use the 5 in it

OpenStudy (anonymous):

is it y=x^2-5 all over 2?

OpenStudy (jdoe0001):

yeap, so \(\bf f(x)={\color{red}{ y}}=\sqrt{2{\color{blue}{ x}}+5}\qquad inverse\implies {\color{blue}{ x}}=\sqrt{2{\color{red}{ y}}+5} \\ \quad \\ \cfrac{{\color{blue}{ x}}^2-5}{2}=y=f^{-1}({\color{blue}{ x}})\qquad f^{-1}({\color{blue}{ 5}})=\cfrac{{\color{blue}{ 5}}^2-5}{2}\)

OpenStudy (anonymous):

10

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

yay, do u mind helping me with another question?

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

write a quadratic inequality for which the solution is -2<x<5